Shift report 20000723 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Jul 23 2000 - 00:04:56 EDT

  • Next message: http account: "Shift report 20000723 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Jul 23 00:04:56 2000, Shift: 16:00-24:00
    Supervisor: YKL, with: TKO
    Log book#3 pages 47 - 47
    
    Shift Summary:
    ------------------------------------------------------------
    RHIC continued ramp tuning most of the evening. There was
    a brief success in ramping but an attempt to cog lost
    blue beam.
    
    HV control has still problems connecting to H1,H2,C1,TOFW.
    KO came at midnight and is trying to remedy it.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Jul 23 2000 - 01:02:55 EDT